# Frequency-Aware Gaussian Splatting Decomposition

### v1

@@ 4.3 Progressive rendering
Similar to other Level-Of-Detail methods, our method naturally supports progressive rendering, enabling adaptive rendering strategies based on computational resources. On edge devices, such as smartphones and laptops, one can render only the low-frequency Gaussians, either at reduced resolution or at full resolution with lower detail, achieving high frame rates with minimal computational cost. As more powerful GPUs become available, additional high-frequency Gaussians can be rendered to progressively refine the image, as shown in figure 4.
Figure 5: Progressive Rendering Consistency: Comparing between OCTree-GS (top) and our method (bottom) across different frequency levels (left to right). In OCTree-GS, lower levels suffer from missing geometric details, which are recovered as more levels are added, leading to artifacts and inconsistent high-frequency details in lower levels. In contrast, our representation preserves the scene’s geometry across all levels, with fine details progressively added, making it more suitable for frequency-aware applications.
Furthermore, our approach is well-suited for streaming applications, where only the low-frequency Gaussians are to be transmitted first, allowing the end user to start rendering immediately. As demonstrated in Figure 4, transmitting more Gaussians further improves reconstruction quality. See further examples in the supplemental.
Figure 5 compares our approach with that of OCTree-GS [12]. As can be seen, our approach maintains a visually coherent approximation of the scene, as opposed to OCTree-GS.
@@ 4.4 Faster Geometry Interaction
Figure 6: Applications of Frequency-Based Gaussian Rendering: (a) Efficient geometry interaction using low-frequency levels for faster processing. Closer points are brighter. Using only low frequency Gaussians (left) is more than twice faster than using all Gaussians (right) with negligible error. (b) Foveated rendering, improving FPS performance using eye-tracking. (c) Frequency-based artistic effects, demonstrating an X-ray-like visualization. Please zoom in for details.
One important application of our frequency decomposition method is fast geometry interaction. Since low-frequency Gaussians define the general structure of the scene, many geometric operations can be efficiently approximated by operating only on this subset. See Figure 5.
We compare the performance of nearest-neighbor search on a randomly sampled 100K point cloud with a scale of 1, using either the full Gaussian set or only the low-frequency subset. Querying the KDTree [2] built from the full model required 1.95 seconds, whereas querying the KDTree constructed from only the low-frequency Gaussians reduced this to 0.88 seconds, achieving a 2.21×2.21$\times$2.21 × speedup. Despite this simplification, the median discrepancy was $0.0006$, and the mean discrepancy was only $0.0018$, corresponding to a $0.18\%$ error relative to the point cloud scale. This result suggests that many geometry-based interactions, such as collision detection and spatial querying, can be significantly accelerated using only the low-frequency representation, with minimal approximation error. The point cloud and the measured scene can be seen in Fig. 6(a).
@@ 4.6 Artistic 3D Filters
Our representation also enables consistent 3D artistic filtering by selectively modifying the attributes of Gaussians at different frequency levels. By altering specific groups of Gaussians, we can achieve artistic effects that remain coherent across different viewpoints. To demonstrate that, one can use our representation to create X-ray style filter, where low-frequency Gaussians are assigned a constant color of dark blue to simulate an underlying structure, while high-frequency Gaussians have their RGB values increased by 0.2, enhancing the visibility of fine details. As demonstrated in figure 6(c), this manipulation creates a 3D-consistent X-ray effect, preserving structural integrity across views. Additional artistic effects on different scenes are showcased in the supplementary material.
Figure 7: Frequency Aware 3D Editing: Our method selectively renders only low-frequency gaussians in the scene, except for gaussians which belongs to the prompted object. The top scene got the prompt ”Lego”, while the scene at the bottom got ”Desk Mat”.
Dataset PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # Gauss $\downarrow$
Ref Ours Ref Ours Ref Ours Ref Ours
Mip-NeRF360 29.763 29.651 0.875 0.874 0.146 0.156 824474 789182
Tanks & Temples 24.752 24.399 0.883 0.879 0.102 0.120 480590 493681
Deep Blending 30.288 30.304 0.912 0.914 0.160 0.173 863849 707819
Table 1: Mean reconstruction metrics (PSNR, SSIM, LPIPS) for Mip-NeRF360, Tanks & Temples, and Deep Blending datasets, comparing SplatFacto (Ref) and our model (Ours). The table also includes the average number of Gaussians used in the reconstruction. For this comparison, we rendered all Gaussian levels in our model. All scenes were trained with 3 frequency levels. SplatFacto performs slightly better than our approach. Visually, the differences are barely noticeable.
# Levels PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
1 (SplatFacto) 33.683 0.957 0.042
2 33.539 0.957 0.045
3 33.499 0.955 0.049
4 33.289 0.952 0.052
Table 2: Trade-off between the number of frequency groups and reconstruction accuracy. Increasing the number of frequency levels slightly reduces reconstruction accuracy.
4.7 Frequency-Aware 3D Editing
Our method enables 3D-consistent, frequency-aware editing by grouping Gaussians into interpretable frequency sub-bands. This allows selective edits while preserving scene consistency across multiple views, overcoming the limitations of 2D masks, which can introduce segmentation artifacts and fail to maintain multi-view coherence.
We leverage Lang-Segment-Anything [9], which extends SAM2 [11] with text-promptable segmentation. Given a user-defined prompt, SAM2 generates object masks for $N$ randomly selected training images. To identify Gaussians corresponding to the prompted object, we project all Gaussians centers onto each selected image and compute the ratio of occurrences inside a mask relative to successful projections. We preserve Gaussians that either belong to the lowest frequency level or have a mask-projection ratio above a certain threshold, resulting in object-focused rendering where the rest of the scene remains blurred. Importantly, this method not only improves segmentation reliability but also enhances efficiency by rendering fewer Gaussians instead of rendering the entire scene and applying post-processing blurring. Furthermore, this enables the storage of smaller scene representations, allowing emphasis on specific objects without requiring the full dataset.
The entire process runs in seconds to a few minutes, depending on the size of the scene and the number of selected images. Figure 7 illustrates frequency-aware object focus across different text prompts, with additional qualitative results in the supplementary material.
@@ 4.8 Ablation Studies
4.9 Limitations
While our method focuses on improving various tasks by leveraging frequency-aware decomposition, it comes with some slight trade-offs and limitations.
We compare our approach to standard SplatFacto 3D-GS and find there is slight degradation in quality, as can be seen in Table 1. Visually, the differences were hardly noticable.
Table 2 analyzes the impact of increasing the number of frequency groups on reconstruction accuracy for the ”kitchen” scene from the Mip-NeRF360 dataset. As the number of frequency levels increases, PSNR and SSIM exhibit a slight decline, while LPIPS increases. This trend suggests that although higher-frequency decomposition enhances flexibility, it introduces minor degradation in reconstruction quality. Consequently, determining the optimal number of frequency levels requires balancing high-fidelity reconstruction with the benefits of frequency separation.
An additional limitation of our method is the increased training time, which results from the need to regulate multiple frequency levels in both the image and frequency domains. This added complexity leads to longer optimization times compared to the baseline model. However, there is no impact on rendering time or memory consumption during inference, ensuring that our method remains efficient for real-time applications.
@@ 5 Conclusions
We introduced frequency decomposition into 3D Gaussian Splatting, inspired by Laplacian pyramids in image processing, to provide structured control over frequency components in a scene. By organizing Gaussians into hierarchical subbands, incorporating Residual Color Gaussians, and leveraging progressive training, our method enhances interpretability and enables efficient level-of-detail management. This approach paves the way for applications such as frequency-aware 3D editing and artistic filtering, allowing precise manipulation of frequency bands for creative effects and content-aware modifications. Furthermore, it supports practical scenarios like foveated rendering, progressive rendering, streaming, and fast geometry interaction, making real-time performance more adaptable across different devices. These capabilities establish our method as a versatile and powerful extension to standard 3D Gaussian Splatting.
@@ 6 Supplementary Material
  • 6.1 Analysis of Gaussian Scale Distribution Across Levels: Distribution analysis of Gaussian scales across frequency levels.
  • 6.2 Foveated Rendering: Implementation details of our foveated rendering approach.
  • 6.3 Comparison to 3D-GS LOD Methods: Evaluation of our method against multi-scale 3D Gaussian Splatting baselines.
  • 6.4 Efficient Multi-Resolution Rendering with Frequency-Decomposed Gaussians: Quality analysis of our hierarchical Gaussian decomposition for rendering at low resolutions.
  • 6.5 Ablation Studies: Detailed discussion on key design choices, including residual color Gaussians, frequency magnitude loss, and image loss at lower levels.
  • 6.6 Additional Examples of Level of Detail: Visualization of level-of-detail decomposition in various scenes.
  • 6.7 Additional Examples of Foveated Rendering: Demonstrations of foveated rendering applied to different scenes.
  • 6.8 Additional Examples of Frequency-Aware 3D Editing: Illustration of selective object emphasis using our frequency-aware editing approach in different scenes.
  • 6.9 Additional 3D Artistic Filters: Examples of artistic filters applied to different scenes using our method.
@@ 6.2 Foveated Rendering
Method Mip-NeRF360 Tanks & Temples Deep Blending
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Multi-scale 3DGS 27.52 - 0.142 23.74 - 0.096 29.70 - 0.096
OCTree-GS 28.05 0.819 0.214 24.68 0.866 0.153 30.49 0.912 0.241
SplatFacto 29.763 0.875 0.146 24.752 0.883 0.102 30.288 0.912 0.160
Ours 29.651 0.874 0.156 24.399 0.879 0.120 30.304 0.914 0.173
Table 3: Comparison of Multi-scale 3D Gaussian Splatting, OCTree-GS, SplatFacto, and Our Method in terms of PSNR, SSIM, and LPIPS. Higher PSNR and SSIM are better, lower LPIPS is better.
Model MipNeRF-360 Tanks & Temples Deep Blending
x4 x2 x4 x2 x4 x2
Splatfacto 27.739 29.536 23.577 24.819 29.446 30.545
Ours 28.443 29.716 23.812 24.526 29.733 30.597
Table 4: PSNR Comparison between Splatfacto (Reference) and Our Method across different datasets and downscale factors.
6.3 Comparison to 3D-GS LOD Methods
In Table 3, we provide a quantitative comparison of our method against Multi-scale 3D Gaussian Splatting (3DGS) [17], OCTree-GS [12], and SplatFacto [16] in terms of PSNR, SSIM, and LPIPS on Mip-NeRF 360 [1], Tanks & Temples [6] and Deep Blending [4] datasets. Since our approach is implemented in a different framework [15] than Multi-scale 3DGS and OCTree-GS [5], certain implementation details, including downscale resolutions, specific hyperparameter settings and optimization strategies, may not be fully identical. Despite these differences, we include their reported results as a point of reference to provide broader context for our performance evaluation. Our method achieves competitive results to prominent 3D-GS LOD methods, demonstrating its effectiveness in high-fidelity scene reconstruction.
6.4 Efficient Multi-Resolution Rendering with Frequency-decomposed Gaussians
In our method, we use the hierarchical structure of our Gaussian representation to render at lower resolutions efficiently. Specifically, with a 3-level trained model, for x4 downscaled rendering, we use only the first level of our hierarchy, while for x2 downscale, we utilize the first two levels. In contrast, the reference method (Splatfacto) directly renders at the target resolution without such hierarchical decomposition. As shown in Table 4, our approach consistently achieves higher PSNR values across different datasets and downscale factors, demonstrating its superior rendering quality at low resolutions. Additionally, since our method processes only a subset of the Gaussians for lower resolutions, it reduces rendering time while maintaining high fidelity.
@@ 6.5 Ablation Studies
Effect of Residual Color Gaussians. We evaluated the effectiveness of Residual Color Gaussians by training a model where all Gaussians retained colors in the standard $[0,1]$ range, rather than using the proposed $[-1,1]$ residual representation for higher-level Gaussians. This modification led to a significant increase in the total number of Gaussians, from 409K to 515K, representing a 20% growth in scene size. While this increase contributed to a slight improvement in PSNR (+0.2 dB), it also resulted in higher memory consumption and slower rendering times. These findings suggest that the residual color representation enables more compact and efficient scene encoding by allowing individual Gaussians to contribute more effectively to high-frequency details. The ability to add or remove color at different hierarchical levels grants each Gaussian greater expressive power, reducing the overall number of Gaussians required for accurate reconstruction.
Effect of Image Loss on Lower Frequency Levels. To assess the importance of applying the image loss at lower frequency levels, we trained a three-level model while omitting the image loss at lower levels. The resulting model achieved a comparable PSNR to our full method when evaluating the final rendered image; however, its intermediate frequency levels exhibited significant degradation. Specifically, when rendering downscaled resolutions using only the first level (corresponding to a ×$\times 4$ downscaled output) or the first two levels (×$\times 2$ downscaled output), we observed a substantial loss of detail, leading to smeared and blurry reconstructions. Quantitatively, our full model achieved a PSNR of 29.65 dB for ×$\times 4$ downscaled rendering and 32.27 dB for ×$\times 2$ downscaled rendering. In contrast, the ablated model, trained without image loss at lower frequency levels, suffered a notable performance drop, with PSNR values of 25.97 dB and 29.36 dB for ×$\times 4$ and ×$\times 2$ downscaled renderings, respectively. These results highlight the crucial role of enforcing spatial consistency across frequency levels, ensuring that intermediate levels remain faithful to their expected resolution.
@@ 6.7 Additional Examples of Foveated Rendering
6.8 Additional Examples of Frequency-Aware 3D Editing
To further demonstrate the effectiveness of our frequency-aware 3D editing approach, we provide additional examples in Figure 12. These results highlight how our method enables selective object emphasis while maintaining 3D consistency. Each row in the figure presents different objects highlighted from a scene using text-based prompts, showcasing the robustness of our 3D voting mechanism. The approach effectively suppresses background details while preserving high-frequency structures for the selected objects, enabling seamless editing without introducing segmentation artifacts.
Figure 12: Additional results of frequency-aware 3D editing. Each row shows a different object selected using a text prompt, demonstrating our method’s ability to produce consistent 3D object emphasis.

### v2

@@ 4.3 Progressive Rendering
Figure 4: Progressive Rendering Consistency: Comparing between OCTree-GS (top) and our method (bottom) across different frequency levels (left to right). In OCTree-GS, lower levels suffer from missing geometric details, which are recovered as more levels are added. In contrast, our representation preserves the scene’s geometry across all levels, with fine details progressively added, making it more suitable for frequency-aware applications.
Progressive rendering on resource-constrained devices allows users to render only low-frequency levels for high frame rates, then progressively add higher-frequency details as computational budget allows. For streaming applications, low-frequency Gaussians can be transmitted first for immediate rendering, with higher-frequency content streamed progressively to refine quality over time.
Our frequency-based decomposition naturally enables progressive rendering where each level provides a meaningful, artifact-free approximation of the scene. Unlike spatial LOD methods that often introduce noticeable artifacts (Figure 4), our frequency hierarchy ensures smooth quality transitions and coherent approximations at all levels.
As shown in Figure 5, our method consistently achieves the highest quality (PSNR) across all LODs while surpassing baselines in rendering speed from Level 3 onward.
Figure 5: Progressive Rendering and Streaming: Comparison of our method, FLOD, and OCTree-GS across different levels of detail. The plot shows the trade-off between rendering speed (FPS) and reconstruction quality (PSNR), where each point corresponds to a specific LOD level and bubble size reflects the number of Gaussians. Results are averaged over all Mip-NeRF 360 scenes. Our method demonstrating superior performance in both quality and speed in the highest three LODs.
@@ 4.4 Faster Geometry Interaction
Figure 6: Applications of Frequency-Based Gaussian Rendering: (a) Efficient geometry interaction using low-frequency levels for faster processing. Closer points are brighter. Using only low frequency Gaussians (left) is more than twice as fast as using all Gaussians (right) with negligible error. (b) Foveated rendering, improving FPS performance using eye-tracking. (c) Frequency-based artistic effects, demonstrating an X-ray-like visualization. Please zoom in for details.
One important application of our frequency decomposition method is fast geometry interaction. Since low-frequency Gaussians define the general structure of the scene, many geometric operations can be efficiently approximated by operating only on this subset. Figure 4 shows that we have structural integrity in lower levels.
We compare the performance of nearest-neighbor search on a randomly sampled 100K point cloud with a scale of 1, using either the full Gaussian set or only the low-frequency subset. Querying the KDTree [4] built from the full model required 1.95 seconds, whereas querying the KDTree constructed from only the low-frequency Gaussians reduced this to 0.88 seconds, achieving a $2.21\times$ speedup. Despite this simplification, the median discrepancy with respect to the KDTree operating on the full model was $0.0006$, and the mean discrepancy was only $0.0018$, corresponding to a $0.18\%$ error relative to the point cloud scale. This result suggests that many geometry-based interactions, such as collision detection and spatial querying, can be significantly accelerated using only the low-frequency representation, with minimal approximation error. The point cloud and the measured scene can be seen in Fig. 6(a).
@@ 4.5 Foveated Rendering
4.6 Promptable 3D Focus
Figure 7 demonstrates our text-prompted frequency-aware 3D focus, where users specify objects via natural language to achieve object-focused rendering that preserves high-frequency detail for targets while blurring the background. This overcomes limitations of 2D mask-based methods, that suffer from artifacts and poor multi-view consistency.
We leverage Lang-Segment-Anything [18] for object identification. Given a text prompt, we generate masks across N training images, then identify corresponding Gaussians by projecting centers and computing mask overlap ratios. We preserve Gaussians that either (1) belong to the base frequency level or (2) exceed the mask threshold, producing 3D-consistent views with sharp objects and blurred backgrounds.
This approach offers both visual and efficiency gains: fewer Gaussians are rendered, enabling compact object-focused representations.
Figure 7: Promptable 3D Focus: Our method selectively renders only low-frequency Gaussians in the scene, except for Gaussians that belong to the prompted object. The left scene got the prompt ”Lego”, while the right scene got ”Desk Mat”.
@@ 4.7 Artistic 3D Filters
Our representation also enables consistent 3D artistic filtering by selectively modifying the attributes of Gaussians at different frequency levels. By altering specific groups of Gaussians, we can achieve artistic effects that remain coherent across different viewpoints. To demonstrate that, one can use our representation to create an X-ray-style filter, where low-frequency Gaussians are assigned a constant color of dark blue to simulate an underlying structure, while high-frequency Gaussians have their RGB values increased by 0.2, enhancing the visibility of fine details. As demonstrated in Figure 6(c), this manipulation creates a 3D-consistent X-ray effect, preserving structural integrity across views. Additional artistic effects on different scenes are showcased in the supplementary material.
@@ 5 Conclusions
We introduce frequency decomposition into 3D Gaussian Splatting, inspired by Laplacian pyramids in image processing, to provide structured control over frequency components in a scene. By organizing Gaussians into hierarchical subbands, incorporating Residual Color Gaussians, and leveraging progressive training, our method enhances interpretability and enables efficient level-of-detail management.
Quantitatively, our method achieves state-of-the-art reconstruction quality and rendering speed among all level-of-detail extensions of 3D Gaussian Splatting, as demonstrated on standard benchmarks. Additionally, this approach paves the way for applications such as promptable 3D focus and artistic filtering, allowing precise manipulation of frequency bands for creative effects and content-aware modifications. Furthermore, it supports practical scenarios like foveated rendering, progressive rendering, streaming, and fast geometry interaction, making real-time performance more adaptable across devices. These capabilities establish our method as both a versatile representation and a strong baseline for frequency-aware 3D scene modeling.
Supplementary Material
@@ 1 Supplementary Material
  • 1.1 Analysis of Gaussian Scale Distribution Across Levels: Distribution analysis of Gaussian scales across frequency levels.
  • 1.2 Foveated Rendering: Implementation details of our foveated rendering approach.
  • 1.3 Ablation Studies: Detailed discussion on key design choices, including residual color Gaussians, frequency magnitude loss, and image loss at lower levels.
  • 1.4 Gaussian Count Dynamics: Analysis of how the number of Gaussians evolves throughout training and across frequency levels.
  • 1.5 Additional Examples of Level of Detail: Visualization of level-of-detail decomposition in various scenes.
  • 1.6 Additional Examples of Foveated Rendering: Demonstrations of foveated rendering applied to different scenes.
  • 1.7 Additional Examples of Promptable 3D Focus: Illustration of selective object emphasis using our promptable 3D focus application in different scenes.
  • 1.8 Additional 3D Artistic Filters: Examples of artistic filters applied to different scenes using our method.
@@ 1.3 Ablation Studies
Effect of Residual Color Gaussians. We evaluated the effectiveness of Residual Color Gaussians by training a model where all Gaussians retained colors in the standard $[0,1]$ range, rather than using the proposed [1,1][-1,1] residual representation for higher-level Gaussians. This modification led to a significant increase in the total number of Gaussians, from 409K to 515K, representing a 20% growth in scene size. While this increase contributed to a slight improvement in PSNR (+0.2 dB), it also resulted in higher memory consumption and slower rendering times. These findings suggest that the residual color representation enables more compact and efficient scene encoding by allowing individual Gaussians to contribute more effectively to high-frequency details. The ability to add or remove color at different hierarchical levels grants each Gaussian greater expressive power, reducing the overall number of Gaussians required for accurate reconstruction.
Effect of Image Loss on Lower Frequency Levels. To assess the importance of applying the image loss at lower frequency levels, we trained a three-level model while omitting the image loss at lower levels. The resulting model achieved a comparable PSNR to our full method when evaluating the final rendered image; however, its intermediate frequency levels exhibited significant degradation. Specifically, when rendering downscaled resolutions using only the first level (corresponding to a $\times 4$ downscaled output) or the first two levels ($\times 2$ downscaled output), we observed a substantial loss of detail, leading to smeared and blurry reconstructions. Quantitatively, our full model achieved a PSNR of 29.65 dB for $\times 4$ downscaled rendering and 32.27 dB for $\times 2$ downscaled rendering. In contrast, the ablated model, trained without image loss at lower frequency levels, suffered a notable performance drop, with PSNR values of 25.97 dB and 29.36 dB for $\times 4$ and $\times 2$ downscaled renderings, respectively. These results highlight the crucial role of enforcing spatial consistency across frequency levels, ensuring that intermediate levels remain faithful to their expected resolution.
@@
1.4 Gaussian Count Dynamics
Figure 10: Evolution of the number of Gaussians during training. Each new frequency level increases the number of Gaussians, but adaptive density control quickly prunes low-utility ones, so the model converges to a compact, lightweight representation.
Figure 10 illustrates the evolution of the number of Gaussians during training. Although each new frequency level initially doubles the number of Gaussians, the adaptive density control remains active throughout the entire training process, continuously removing low-utility components. As a result, the total number of Gaussians quickly stabilizes after each addition, resulting in a compact and efficient representation. This dynamic ensures that the final model remains lightweight despite the hierarchical structure.
@@ 1.6 Additional Examples of Foveated Rendering
1.7 Additional Examples of Promptable 3D Focus
To further demonstrate the effectiveness of our promptable 3D focus application, we provide additional examples in Figure 13. These results highlight how our method enables selective object emphasis while maintaining 3D consistency. Each row in the figure presents different objects highlighted from a scene using text-based prompts, showcasing the robustness of our 3D voting mechanism. The approach effectively suppresses background details while preserving high-frequency structures for the selected objects, enabling seamless editing without introducing segmentation artifacts.
Figure 13: Additional results of our promptable 3D focus application. Each row shows a different object selected using a text prompt, demonstrating our method’s ability to produce consistent 3D object emphasis.